What is the length of a diagonal of a sqaure with a perimeter of 28 centimeters?
Answer by stanbon(75887)

What is the length of a diagonal of a square
with a perimeter of 28 centimeters?
---
Perimeter of a square = 4s where s is the length of each side.
------------------
4s = 28
s = 7 cm (length of each side)
----
Note: diagonal^2 = side^2 + side^2
-----------------
d^2 = 2s^2
d^2 = 2*49
d = 7*sqrt(2) cm
==================
Cheers,
Stan H.
========